# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ 6.3.2 Unobserved Treatment Heterogeneity
Now, we show how our assumptions can be modified to accommodate some forms of unobserved treatment heterogeneity. Instead of viewing policy effects as a function of observed covariates $X$ only, we now model the policy effects at the site level. Specifically, for each $s\in\mathcal{S}$, denote by $\tau_{s}\in\mathbb{R}$ the policy effect in site $s$. Then, instead of viewing $\tau$ as a function, we will let $\tau:=(\tau_{1},\tau_{2},...,\tau_{S})^{\top}\in\mathbb{R}^{S}$ be a vector of dimension $S$ that represents the policy effects of all sites in $\mathcal{S}$. Then, Assumptions 1 and 2 can be replaced with the following:
@@
For any $s,s^{\prime}\in\mathcal{S}$, $s\neq s^{\prime}$, $\tau$ satisfies:
@@
Assumption 4 allows sites with the same covariates to have different policy effects. The difference, however, is assumed to be at most $c$. We motivate Assumption 4 using a simple linear regression model in Appendix B.4. Denote by $\mathcal{F}_{C,c}\subset\mathbb{R}^{S}$ the collection of all vectors of dimension $\mathbb{R}^{S}$ satisfying (36). For this parameter class, the minimax-regret optimal purposive sampling scheme and treatment rule are redefined to solve
$\inf_{\mathscr{S}\in\mathcal{A}(k),T\in\mathcal{T}_{\mathscr{S}}^{1/2}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau).$ (37)
It turns out that the purposive sampling scheme that solves (37) can still be approximated by the solution of the $k$-median problem in Definition 3 with a different cost function.
To see this, note that the functional class $\mathcal{F}_{C,c}$ is still convex and centrosymmetric. Thus, we can show that, for the evidence aggregation problem studied in Montiel Olea et al. (2023), the conclusion of their Proposition 1(iii) extends to $\mathcal{F}_{C,c}$ as well. Then, under Assumption 4, Lemmas 1 and 2 continue to hold with minor modifications. In particular, when $C$ is large enough (with a threshold that can be exactly characterized), we have that
$\inf_{T\in\mathcal{T}_{\mathscr{S}}^{1/2}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau)\leq\;\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}\right)+\left(\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}\right).$
In addition, the following lower bound can also be verified to hold
$\inf_{T\in\mathcal{T}_{\mathscr{S}}^{1/2}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\left(\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}\right).$
The two bounds above imply that, even when treatment effect heterogeneity is characterized by Assumption 4, the optimized value of the minimax problem (37) can still be uniformly approximated by the following objective function
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\left(\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}\right)\right).$
@@
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\left(\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}\right)\right).$
This problem is equivalent to the $k$-median problem in Definition 3, where the connection cost between sites $s$ and $s^{\prime}$ is given by $m(X_{s},X_{s^{\prime}})$ (neither $C$ nor $c$ are required as inputs).
@@ 7 Conclusion
This paper presented a decision-theoretic justification for viewing the question of how to best choose where to experiment in order to optimize external validity as a $k$-median (clustering) problem. More concretely, we presented conditions under which minimizing the worst-case, welfare-based regret among all purposive (nonrandomized) schemes that select $k$ sites is approximately equal, and can be exactly equal, to finding the $k$ most central vectors of baseline site-level covariates.
We believe there are many interesting directions for future work. For example, while we focused on purposive sampling schemes, it would be interesting to better understand the value of randomized sampling schemes and whether site-level covariates can be used to design such randomized selection with an eye on external validity. We also think that discussions around the relation between the $k$-median problem and synthetic purposive sampling of Egami and Lee (2024) open interesting research directions to provide a decision-theoretic justification for the use the synthetic control of Abadie et al. (2010). Finally, it would be interesting to explore the use of (mixed) integer programming techniques to solve for the purposive sampling schemes suggested in Egami and Lee (2024) and Gechter et al. (2024).
@@ A.1 Proof of Lemma 1
Fix the selected sites $\mathscr{S}$, and denote the cardinality of $\mathscr{S}$ as $\textrm{card}(\mathscr{S})$. Let $\mathscr{S}_{1}<\mathscr{S}_{2}<\ldots<\mathscr{S}_{\textrm{card}(\mathscr{S})}$ denote the indices of the $\textrm{card}(\mathscr{S})$ experimental sites in $\mathscr{S}$. For a given experimental site $s\in\mathscr{S}$, let $\widehat{\mathcal{\tau}}_{s}$ denote its corresponding treatment effect estimate. Let
$\widehat{\tau}_{\mathscr{S}}:=(\hat{\tau}_{\mathscr{S}_{1}},\ldots,\hat{\tau}_{\mathscr{S}_{\textrm{card}(\mathscr{S})}})^{\top}$
@@
Consider the decision rule, $T^{*}\in\mathcal{T}^{1/2}_{\mathscr{S}}$, that recommends, for each policy-relevant site $s\in\mathcal{S}_{P}$ and given data $\widehat{\tau}_{\mathscr{S}}$, the following action:
@@
    $\tilde{\sigma}_{s}:=\sqrt{\left(\frac{C\|X_{s}-X_{N_{\mathscr{S}}(s)}\|}{\sqrt{\pi/2}}\right)^{2}-\sigma^{2}_{s}}.$
Note that the expression in ii) above is well-defined for every $s\in\mathcal{S}_{P}$ when $C$ is large enough. This decision rule is the minimax-regret optimal rule (provided $C$ is large enough) for the evidence aggregation framework discussed in Yata (2021) and Montiel Olea et al. (2023), both of which build upon Stoye (2012), and is indeed in $\mathcal{T}^{1/2}_{\mathscr{S}}$.
Define $C(\mathscr{S})$ to be any value of $C$ for which $\tilde{\sigma}_{s}>0$ for every $s\in\mathcal{S}_{P}$. Note that by definition of infimum,
@@
Moreover, it follows from (Montiel Olea et al., 2023, Proposition 1 (iii) and its proof) that if $C>\underset{s\in\mathcal{S}_{P}\backslash\mathscr{S}}{\max}\left\{\sqrt{\frac{\pi}{2}}\frac{\sigma_{N_{{\mathscr{S}}(s)}}}{\left\|X_{s}-X_{N_{{\mathscr{S}}(s)}}\right\|}\right\}:=C(\mathscr{S})$, where $\sigma_{N_{{\mathscr{S}}(s)}}$ denotes the standard deviation of the nearest neighbor estimate in $\mathscr{S}$ for site $s\in\mathcal{S}_{P}\backslash\mathscr{S}$, it holds that for any $s\in\mathcal{S}_{P}\backslash\mathscr{S}$:
@@
$\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\leq\;\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}\right)+\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$ (40)
@@ A.2 Proof of Lemma 2
[…]
@@
We start by showing that the function $\tau^{*}$ is Lipschitz with constant $C$. To see this, consider three cases:
Case 1: Suppose first that $X,X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}\}$. In this case, we trivially have
@@
Case 2: Suppose now that $X\notin\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}\}$, but $X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}\}$. In this case,
@@
where the last inequality follows by the definition of $X_{\mathcal{N}_{\mathscr{S}}}$ and the fact that $X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}\}$.
Case 3: Finally, take $X,X^{\prime}\notin\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}\}$. In this case,
$|\tau^{*}(X)-\tau^{*}(X^{\prime})|=C|\|X-X_{\mathcal{N}_{\mathscr{S}}}\|-\|X^{\prime}-X^{\prime}_{\mathcal{N}_{\mathscr{S}}}\||.$
@@
Part II: Since $\tau^{*}\in\textrm{Lip}_{C}(\mathbb{R}^{d})$, then for any treatment rule $T$:
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\mathcal{R}(T,\mathscr{S},\tau^{*}).$ (41)
Moreover, by definition,
@@
$\mathcal{R}(T,\mathscr{S},\tau^{*})$ $=$ $\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\tau^{*}(X_{s})\left(\mathbf{1}\{\tau^{*}(X_{s})\geq 0\}-\mathbb{E}_{\tau^{*}_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right),$ (42)
@@
where the last line uses the definition of $\tau^{*}$. Equations (41) and (42) thus imply that for any $T\in\mathcal{T}^{1/2}_{\mathscr{S}}$:
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$ (43)
@@
$\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$

### v2

@@ 6.3.2 Unobserved Treatment Heterogeneity
Our assumptions can also be modified to accommodate some forms of unobserved treatment heterogeneity. Instead of viewing policy effects as a function of observed covariates $X$ only, we now model the policy effects at the site level. Specifically, for each $s\in\mathcal{S}$, denote by $\tau_{s}\in\mathbb{R}$ the policy effect in site $s$. Then, instead of viewing $\tau$ as a function, we will let $\tau:=(\tau_{1},\tau_{2},...,\tau_{S})^{\top}\in\mathbb{R}^{S}$ be a vector of dimension $S$ that represents the policy effects of all sites in $\mathcal{S}$. Then, Assumptions 1 and 2 can be replaced with the following:
@@
For any $^{\prime}\in\mathcal{S}, sss\neq s^{\prime}$, $\tau$ satisfies:
@@
Assumption 4 allows sites with the same covariates to have different policy effects. The difference, however, is assumed to be at most $c$. We motivate Assumption 4 using a simple linear regression model in Appendix B.4. Denote by $\mathcal{F}_{C,c}\subset\mathbb{R}^{S}$ the collection of all vectors of dimension $\mathbb{R}^{S}$ satisfying (16). For this parameter class, the MMR purposive sampling scheme and treatment rule are redefined to solve
$\inf_{\mathscr{S}\in\mathcal{A}(k),T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau).$ (17)
Then the purposive sampling scheme that solves (17) can still be approximated by the solution of the $k$-median problem in Definition 3 with a different cost function. To see this, note that the functional class $\mathcal{F}_{C,c}$ is still convex and centrosymmetric. Thus, we can show that, for the evidence aggregation problem studied in Montiel Olea et al. (2025), the conclusion of their Proposition 1(iii) extends to $\mathcal{F}_{C,c}$ as well. Then, under Assumption 4, Lemmas 1 and 2 continue to hold with minor modifications. In particular, when $C$ is large enough (with a threshold that can be exactly characterized), we have that
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau)~\leq~\;\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}~+~\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}.$
In addition, the following lower bound can also be verified to hold:
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\mathcal{F}_{C,c}}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}.$
These bounds imply that, even when treatment effect heterogeneity is characterized by Assumption 4, the optimized value of the minimax problem (17) can still be uniformly approximated by
$\inf_{\mathscr{S}\in\mathcal{A}(k)}~\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}.$
@@
$\inf_{\mathscr{S}\in\mathcal{A}(k)}~\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\frac{Cm(X_{s},X_{N_{\mathscr{S}}(s)})+c}{2}.$
This problem is equivalent to the $k$-median problem in Definition 3, where the connection cost between sites $s$ and $s^{\prime}$ is given by $m(X_{s},X_{s^{\prime}})$.
6.3.3 Treatment Heterogeneity Implied by Structural Model
Our approach can also accommodate structural models of treatment heterogeneity. To see this, let $\tau(X_{s})=g(\theta,X_{s})$ be the treatment effect of site $s$, built from a structural model $g(\cdot,\cdot)$ with parameters $\theta\in\Theta$ (Gechter et al., 2024). Then, the site-level treatment effect is governed by $\tau:=\left\{\tau(\cdotp)\mid\tau(\cdotp)=g(\theta,\cdotp),\theta\in\Theta\right\}$. We may define the metric
$m(x,x^{\prime}):=\sup_{\theta\in\Theta}\left|g(\theta,x)-g(\theta,x^{\prime})\right|,$
implying that $\tau(\cdot)$ informed by the structural model is $1$-Lipschitz continuous with respect to the redefined $m(\cdot,\cdot)$ metric. Moreover, suppose we have a plausible structural estimate $\theta^{*}$ derived from previous studies. One may incorporate such information by analogously defining
$m(x,x^{\prime}):=\sup_{\theta\in\Theta^{*}}\left|g(\theta,x)-g(\theta,x^{\prime})\right|,$
where $\Theta^{*}\subset\Theta$ is a subset of parameter values (possibly a singleton) containing $\theta^{*}$.
6.3.4 General Convex and Centrosymmetric Class
Instead of Assumption 1, suppose $\tau(\cdotp)\in L_{cc}$, where $L_{cc}$ is a convex and centrosymmetric (i.e., $\tau(\cdotp)\in L_{cc}$ implies $\tau(\cdotp)\in L_{cc}$) parameter space. It turns out that we can still derive a tractable upper bound for the associated MMR optimal purposive sampling scheme. For example, suppose $\mathcal{S}_{E}\cap\mathcal{S}_{P}=\emptyset$. Given $\mathscr{S}\in\mathcal{A}(k)$, let
$\overline{I}_{s,\mathcal{\mathscr{S}}}(0):=\sup\left\{u\in\mathbb{R}\mid\tau(X_{s})=u,\tau(X_{\mathscr{S}_{j}})=0,j=1,\ldots,\#\mathscr{S},\tau\in L_{cc}\right\}$
be the upper bound of the identified set for the treatment effect in policy relevant site $s\in\mathcal{S}_{P}$ when the treatment effects for each of the experimental sites in $\mathscr{S}$ equals 0. Applying Yata (2021) and Montiel Olea et al. (2025) for general convex and centrosymmetric parameter space, we may derive the following upper bound (following analogous steps of Lemma 1): For every $\mathscr{S}\in\mathcal{A}(k)$, whenever $\max\limits_{\mathscr{S}\in\mathcal{A}(k),s\in\mathcal{S}_{P}}\left\{\overline{I}_{s,\mathcal{\mathscr{S}}}(0)\right\}$ is sufficiently large, we have
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in L_{cc}}\mathcal{R}(T,\mathscr{S},\tau)\leq\left(\frac{1}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\overline{I}_{s,\mathcal{\mathscr{S}}}(0)\right).$ (18)
Therefore, one may still try to solve the site selection problem by minimizing
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\sum_{s\in\mathcal{S}_{P}}\overline{I}_{s,\mathcal{\mathscr{S}}}(0).$ (19)
However, this foregoes two advantages of Lipschitz restrictions. In terms of algorithms, the objective function in (19) does not give a direct connection to a distance metric given the general nature of $\overline{I}_{s,\mathcal{\mathscr{S}}}(0)$, implying additional computational burden; in terms of theory, it may be more challenging to show that the upper bound in (18) is tight, while in the Lipschitz case, we are able to explicitly construct a function in $\textrm{Lip}_{C}(\mathbb{R}^{d})$ under which the upper bound is attained (Lemma 2).
@@ 7 Conclusion
This paper presented a decision-theoretic justification for viewing the question of how to best choose where to experiment in order to optimize external validity as a $k$-median problem. More concretely, we presented conditions under which minimizing the worst-case, welfare-based regret among all purposive (nonrandomized) schemes that select $k$ sites is approximately equal, and can be exactly equal, to finding the $k$ most central vectors of baseline site-level covariates.
We believe there are many interesting directions for future work. For example, while we focused on purposive sampling schemes, it would be interesting to better understand the value of randomized sampling schemes and whether site-level covariates can be used to design such randomized selection with an eye on external validity. We also think that discussions around the relation between the $k$-median problem and synthetic purposive sampling of Egami and Lee (2024) open interesting research directions to provide a decision-theoretic justification for the use the synthetic control of Abadie et al. (2010).
@@ A.1 Proof of Lemma 1
Fix the selected sites $\mathscr{S}$, and denote the cardinality of $\mathscr{S}$ as $\mathscr{S}$. Let $\mathscr{S}_{1}<\mathscr{S}_{2}<\ldots<\mathscr{S}_{\#\mathscr{S}}$ denote the indices of the $\mathscr{S}$ experimental sites in $\mathscr{S}$. For a given experimental site $s\in\mathscr{S}$, let $\widehat{\mathcal{\tau}}_{s}$ denote its corresponding treatment effect estimate. Let
$\widehat{\tau}_{\mathscr{S}}:=(\widehat{\tau}_{\mathscr{S}_{1}},\ldots,\widehat{\tau}_{\mathscr{S}_{\#\mathscr{S}}})^{\top}$
@@
Consider the decision rule, $T^{*}\in\mathcal{T}_{\mathscr{S}}$, that recommends, for each policy-relevant site $s\in\mathcal{S}_{P}$ and given data $\widehat{\tau}_{\mathscr{S}}$, the following action:
@@
    $\tilde{\sigma}_{s}:=\sqrt{\left(\frac{C\|X_{s}-X_{N_{\mathscr{S}}(s)}\|}{\sqrt{\pi/2}}\right)^{2}-\sigma^{2}_{N_{\mathscr{S}}(s)}},$
and $\sigma_{N_{{\mathscr{S}}(s)}}$ denotes the standard deviation of the treatment effect estimator corresponding to the nearest-neighbor of site $s$ when considering experimental sites $\mathscr{S}$. Note that the expression in ii) above is well-defined for every $s\in\mathcal{S}_{P}\backslash\mathscr{S}$ when $C$ is large enough. Moreover, this decision rule: (1) is the minimax-regret optimal rule (provided $C$ is large enough) for the evidence aggregation framework discussed in Yata (2021) and Montiel Olea et al. (2025), both of which build upon Stoye (2012), and (2) satisfies the following property: for any diagonal matrix $\Sigma$, and $\forall s\in\mathcal{S}_{P}$, $\mathbb{E}[T_{s}(U)]=1/2$, where $U\sim\mathcal{N}_{\#\mathscr{S}}(\mathbf{0},\Sigma)$. That is, $T^{*}$ is such that the ex-ante probability of implementing the policy is $50\%$ whenever the true treatment effects at the sites experimented on are zero.
Define $C(\mathscr{S})$ to be the smallest value of $C$ for which $\tilde{\sigma}_{s}>0$ for every $s\in\mathcal{S}_{P}$. Note that by definition of infimum,
@@
Moreover, it follows from (Montiel Olea et al., 2025, Proposition 1 (iii) and its proof) that if $C>\underset{s\in\mathcal{S}_{P}\backslash\mathscr{S}}{\max}\left\{\sqrt{\frac{\pi}{2}}\frac{\sigma_{N_{{\mathscr{S}}(s)}}}{\left\|X_{s}-X_{N_{{\mathscr{S}}(s)}}\right\|}\right\}:=C(\mathscr{S})$, it holds that for any $s\in\mathcal{S}_{P}\backslash\mathscr{S}$:
@@
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\leq\;\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}\right)+\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$ (22)
@@ A.2 Proof of Lemma 2
Fix the selected sites $\mathscr{S}$ and let $\mathscr{S}_{1}<\mathscr{S}_{2}<\ldots<\mathscr{S}_{\#\mathscr{S}}$ denote the indices of the $\mathscr{S}$ experimental sites in $\mathscr{S}$. For any $X\in\mathbb{R}^{d}$, define $X_{\mathcal{N}_{\mathscr{S}}}$ to be the element in $\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$ that is closest to $X$ in terms of $\|\cdot\|$ (if there is more than one closest element, pick the $X$ associated with the smallest index). The proof has three parts.
@@
We start by showing that $\tau^{*}$ is Lipschitz with constant $C$. To see this, consider three cases:
Case 1: Suppose first that $X,X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$. In this case, we trivially have
@@
Case 2: Suppose now that $X\notin\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$, but $X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$. In this case,
@@
where the last inequality follows by the definition of $X_{\mathcal{N}_{\mathscr{S}}}$ and the fact that $X^{\prime}\in\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$.
Case 3: Finally, take $X,X^{\prime}\notin\{X_{\mathscr{S}_{1}},\ldots,X_{\mathscr{S}_{\#\mathscr{S}}}\}$. In this case,
[…]
@@
Part II: Since $\tau^{*}\in\textrm{Lip}_{C}(\mathbb{R}^{d})$, so is $\tau^{*}$. Consider a choice of $\tau$ such that $\tau=\tau^{*}$ with probability $1/2$ and $\tau=-\tau^{*}$ with probability $1/2$. Then, for any treatment rule $T$:
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\frac{1}{2}\mathcal{R}(T,\mathscr{S},\tau^{*})+\frac{1}{2}\mathcal{R}(T,\mathscr{S},-\tau^{*})$
Moreover, by definition, for all $\tau$,
@@
$\mathcal{R}(T,\mathscr{S},\tau^{*})$ $=$ $\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\tau^{*}(X_{s})\left(\mathbf{1}\{\tau^{*}(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}^{*}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right)$
@@
where the last line uses the definition of $\tau^{*}$. Moreover, we have $\|X_{s}-X_{N_{\mathscr{S}}(s)}\|>0$ for all $s\in\mathcal{S}_{P}\backslash\mathscr{S}$ by Assumption 2, implying $\tau^{*}(X_{s})<0$ for all $s\in\mathcal{S}_{P}\backslash\mathscr{S}$. Therefore, we have, analogously,
$\mathcal{R}(T,\mathscr{S},-\tau^{*})$ $=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}C\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\mathbb{E}_{\mathbf{0}}[T_{s}(\widehat{\tau}_{\mathscr{S}})].$
Conclude that for any $T\in\mathcal{T}$:
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)$ $\geq\frac{1}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}C\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\left(1-\mathbb{E}_{\mathbf{0}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right)$
$+\frac{1}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}C\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\mathbb{E}_{\mathbf{0}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]$
$=\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\left\{1-\mathbb{E}_{\mathbf{0}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]+\mathbb{E}_{\mathbf{0}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right\}$
$=\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|.$ (23)
@@
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|.$